# FlowBotHD: History-Aware Diffuser Handling Ambiguities in Articulated Objects Manipulation

### v1

@@ C.1 Multi-modality of Dataset
For other object categories, there is no obvious automatic way to determine whether an object is ambiguous, it is hard to manually count ambiguous samples for all the categories (and subjective), therefore we analyze conceptually where ambiguities typically happen instead. There are three typical multi-modal objects (See Figure 5) : 1) Joints without handles, 2) Joints with small handles that might be missed when downsampling, and 3) Joints with handles that are still possible to open in different ways: for example a dishwasher/oven with a handle on the top of its front surface, although the handle eliminates many modes, it is unclear whether it opens by pulling out like a drawer, or pulling from the top to the bottom like a lid or door.
@@
Similar to counting modes of general articulated objects, there is no precise way to count multi-modality in predictions without including manual judging and counting. To approximately have an idea about the multi-modal predictions of the model, we make 50 predictions for each object in the test set and calculate their RMSE with respect to the ground truth flow. We record how many of the objects include both RMSE $<$ 0.2 (presumably a correct prediction) and RMSE $>$ 0.6 (presumably a different mode than the ground truth mode). We can see from the bar plot in Figure 6 that the model predicts multi-modality mostly in conceptually multi-modal categories like Door, WashingMachine, and Refrigerators. However it also makes multi-modal predictions with conceptually inambiguous cases, which aligns with limitations of this work we addressed in 7 about the difficult balance between precision and multi-modality.
@@ C.2 Model Performance on Multi-modal Cases
To specifically analyze our model’s improvements on handling multi-modality, we evaluated baseline FlowBot3D trained on randomly opened dataset (RO) / mixed dataset (MD) and our model FlowBotHD on doors. We open each test door from 0% to 100%. We average the metrics within 10% open to compute the performance for closed doors, and we average the metrics above 10% open to compute the performance for open doors. As we can see from Table 4, our model outperforms the baselines by a large margin across all metrics. The observation that FlowBot3D trained on randomly opened dataset outperforms the one trained on the mixed dataset shows that multi-modal training data can confuse the regression models’ training process.
@@ Appendix E Policy Analysis
We conduct further analysis on components of our roll-out policies, including the occasional negative effect of history filter, the efficiency of switch grasp point policy and the effect of step number limitations.
@@ E.2 Switch Grasp Point Analysis
We made 2 supplementary plots to visualize the frequency of switching grasp points in our simulation experiments. Figure 10 shows the switch grasp point rate per step. We can see that the switch grasp point rate gets lower among the first few steps - the consistency increases with multi-modality decreased. But it increases slightly at the later stages - possibly meaning that occlusions could happen and instablize the grasp point predictions. Figure 11 shows the distribution of the number of switch grasps across the dataset. We can see that about 60% of objects only require one grasp action (the first step). And for the other 40%, most of them require only $<$5 extra switches of the grasp point.

### v2

@@ C.1 Multi-modality of Dataset
For other object categories, since there is no generic way to determine whether an object is ambiguous, it is hard to manually count ambiguous samples for all the categories (not to mention ambiguity might be subjective), therefore we describe broadly where ambiguities typically happen. There are three typical multi-modal objects (See Figure 6) : 1) Joints without handles, 2) Joints with small handles that might be erased when downsampling the point cloud, and 3) Joints with handles that may afford different opening modes: for example a dishwasher/oven with a handle on the top of its front surface, although the handle eliminates many modes, it is unclear whether it opens by pulling out like a drawer, or pulling from the top to the bottom like a lid or door.
@@
Similar to counting modes of general articulated objects, there is no precise way to count multi-modality in predictions without including manual labeling. To approximately evaluate the multi-modal predictions of the model, we make 50 predictions for each object in the test set and calculate their RMSE with respect to the ground truth flow. We record how many of the objects include both RMSE $<$ 0.2 (presumably a correct prediction) and RMSE $>$ 0.6 (presumably a different mode than the ground truth mode). We can see from the bar plot in Figure 7 that the model predicts multi-modality mostly in conceptually multi-modal categories like Door, WashingMachine, and Refrigerators. However it also makes multi-modal predictions with conceptually unambiguous cases, which aligns with limitations of this work we addressed in 6 about the difficulty of striking a balance between precision and multi-modality.
@@ C.2 Model Performance on Multi-modal Cases
To specifically analyze our model’s improvements on handling multi-modality, we evaluated baseline FlowBot3D trained on randomly opened dataset (RO) / mixed dataset (MD) and our model FlowBotHD on doors. We open each test door from 0% to 100%. We average the metrics within 10% open to compute the performance for closed doors, and we average the metrics above 10% open to compute the performance for open doors. As we can see from Table 4, our model outperforms the baselines by a large margin across all metrics. We observe that FlowBot3D trained on randomly opened dataset outperforms the one trained on the mixed dataset, showing that multi-modal training data can confuse the regression models’ training process.
@@ Appendix E Policy Analysis
We conduct further analysis on components of our roll-out policies, including the occasional negative effect of the history filter, the efficiency of switching the grasp point during policy execution and the effect of limiting the number of steps during execution.
@@ E.2 Switch Grasp Point Analysis
We made 2 supplementary plots to visualize the frequency of switching grasp points in our simulation experiments. Figure 11 shows the switch grasp point rate per step. We can see that the switch grasp point rate gets lower among the first few steps - the consistency increases with multi-modality decreased. But it increases slightly at the later stages - possibly meaning that occlusions could happen and destabilize the grasp point predictions. Figure 12 shows the distribution of the number of switch grasps across the dataset. We can see that about 60% of objects only require one grasp action (the first step). And for the other 40%, most of them require only $<$5 extra switches of the grasp point.